Name: Catherine Caracciolo

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 10 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 8

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 10

Total: 93 (A-)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 9

3. Is free of grammatical mistakes – 10

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences – 10

6. Uses correct punctuation – 10

7. Is free of jargon, cliches, and misused words – 8

8. Cites quotations and/or references correctly – 10

Total: 97 (A)

Comments: Catherine, even though I disagree pretty strongly with many of your conclusions, I think this is a much better piece of writing than your first essay. You have laid out and forcefully defended a possible criticism of Skloot’s book.

The main thing I think you are not appreciating is the way the elements of the book that you object to (the history of the Tuskegee Syphilis Study, the discussion of oral traditions of “Night Doctors,” and the like) are there to provide a context for the fear and mistrust that many African-Americans (including to some extent Henrietta herself, and even more strongly some in Henrietta’s family) have for the medical profession.

You say (I think) that those historical facts don’t have anything to do with Henrietta Lacks or HeLa. But that, in a way, is part of Skloot’s point – medical researchers have tended to put their work ahead of the interests of the patients who supply the “raw material” – the cells in this case. And that attitude also contributes in a way to the abuses (when they happen – fortunately rarely) because of the way it lets the scientists distance themselves from the human beings whose tissues they are working with. Think of Mary Kubiczek’s statement that up until she saw Henrietta’s body during the autopsy, she had never made the connection that the HeLa cells were coming from an actual person.

I also think you are giving some of the doctors a ”free pass” by uncritically accepting some of their statements at face value rather than asking whether those statements might be a bit self-serving. I’m thinking mainly here of Howard Jones’s defence of the treatment Henrietta received. (Yes, I agree, Skloot’s argument there would be stronger if she gave more direct evidence of the studies she is citing – did you look at the notes at the end, though?) If one was inclined to be cynical, one might also see the free treatment of Henrietta’s deviated septum as less than a selfless act of charity. It might also have been a great chance for doctors in training to get some practice. Is that automatically bad? Of course not, but it’s also not as clear-cut as you are implying.

Finally, whether or not using tissue from a patient without their consent is (or was) legal, can’t (and shouldn’t) we ask another question? Namely, shouldn’t we ask whether it is right, or ethical to do that? You seem to assume that the answer is definitely yes if it advances science (sorry if I am misrepresenting you viewpoint). But isn’t this something reasonable people might disagree about?